i need to find the slope of a line without graphing and i need to write the equation i have 3 problems
1. (11,-2) and (-7,5)
2.(9,2) and (15,4)
3.(3,-7) and (13,8)
i need to write the equation as y= mx + or - B
and i need it now like in 30 minutes maybe

Answer by Alan3354(69443)

1. (11,-2) and (-7,5)
Slope, m = diffy/diffx
m = (-2 -5)/(11 -(-7)) = -7/18
Use either point to find b:
-2 = (-7/18)*11 + b = -77/18 + b
b = 41/18
y = mx + b
y = (-7/18)x + 41/18
It's always +b, b can be + or minus
-------------------------------------
2.(9,2) and (15,4)
3.(3,-7) and (13,8)
i need to write the equation as y= mx + or - B
-----------------
You do the other 2. email me to check your work at Moral Loophole@aol.com